# Segregation in Nuclear Stellar Clusters: Rates and Mass Distributions of TDEs, QPEs, Plunges, and EMRIs

### v1

@@ 3.1 TDE rate
The majority of disrupted stars with masses in the range $1\lesssim m\lesssim 3$ originate from the vicinity of $r_{d}(m)$, as defined by Eq. (9), where they dominate the scatterings. In this region, the relaxation timescale (Eq. 8) is
@@
At smaller distances, both their number and the two-body scattering flux decrease as a power-law, and at larger distances their number is exponentially suppressed (Linial & Sari, 2022).
Therefore, using Eqs. (4), (13), and (15), the mass-dependent tidal disruption rate is given by
@@
$\approx 5\times 10^{-5}m^{-4\gamma/5}\left(\frac{M_{BH}}{M_{\rm MW}}\right)^{-1/4}{\rm yr^{-1}}.$
Consequentially, the enhancement factor (defined in Eq. 14) is
@@
Given that $\gamma\approx 5$ for this stellar mass range (Eq. 5), the TDE rate scales roughly as ${\rm d}\Gamma/{\rm d}m\propto m^{-4}$ and the enhancement factor is $\mathcal{Q}_{TDE}\simeq m$. Notably, the TDE rate, as well as the rates of EMRIs and plunges (calculated below), scales with the SMBH mass as $\Gamma\propto M_{BH}^{-1/4}$ (in agreement with previous results, e.g., Hopman & Alexander, 2005; Kochanek, 2016; Yao et al., 2023).
As discussed in §2.4, the presence of a BH population modifies the cusp within $R_{\bullet}$. The transition of massive stars ($m\gtrsim\widetilde{m}$) from being primarily self-scattered to being scattered by BHs results in a sharp increase in the enhancement factor [^3] This increase originates from the shorter relaxation timescale when scattered by the more massive BHs compared to the stars, $\tau_{\rm rlx}(\widetilde{m})/\tau_{\rm rlx}(\widetilde{m}_{\bullet})\approx 3$, as evident from Eqs. (15) and (20). around $R_{\bullet}$, $\delta\mathcal{Q}\approx\left(\widetilde{m}_{\bullet}/\widetilde{m}\right)^{1/2}\approx 3$. Therefore
$\mathcal{Q}_{TDE}(m\gtrsim\widetilde{m})$ $=\mathcal{Q}_{TDE}(\widetilde{m})\delta\mathcal{Q}$ (18)
@@
The enhancement factor of stellar disruptions is illustrated in Fig. (2) and can be summarized as follows:
@@ 3.2 Plunge rate
The stellar-mass BHs are predominantly scattered by the $\widetilde{m}_{\bullet}\approx 30$ BHs near $R_{\bullet}$. Therefore, the relevant scattering timescale (Eq. 8) for the plunge rate is given by
@@
$\approx 10^{-7}\left(\frac{m_{\bullet}}{10}\right)^{-5/2}\left(\frac{M_{BH}}{M_{\rm MW}}\right)^{-1/4}{\rm yr^{-1}}.$
@@
This scaling reflects the flattening of the BH mass spectrum caused by the segregation into the sphere of influence (see §2.4). It applies to both lighter and heavier BHs [^4] For massive stellar-mass BHs, with $m_{\bullet}\gtrsim 125$, The critical radius $R_{c}$ (Eq. 23), which differentiate between plunges and EMRIs, exceeds $R_{\bullet}$ (Eq. 11), the radius around which the BHs concentrate. Therefore, such BHs are more likely to form EMRIs than undergo a plunge., as they are mostly scattered by BHs with mass $\widetilde{m}_{\bullet}$ around $R_{\bullet}$, leading to $\left.{\rm d}\Gamma/{\rm d}m\right|_{Plunge}\propto m_{\bullet}^{-\widetilde{\gamma}_{\bullet}}$.
@@ 3.3 EMRI rate
$\frac{R_{c}}{R_{h}}$ $\approx\left(\frac{m_{\bullet}}{16\sqrt{2}\log\Lambda\widetilde{m}_{\bullet}^{1/2}}\right)^{4/5}\approx 0.02\left(\frac{m_{\bullet}}{10}\right)^{4/5}.$ (23)
@@
$\approx 5\times 10^{-7}m_{\bullet}^{1-\widetilde{\gamma}_{\bullet}}\left(\frac{125}{m_{\bullet}}\right)^{m_{\bullet}/150}$
×$\quad\times\left(\frac{M_{BH}}{M_{\rm MW}}\right)^{-1/4}{\rm yr^{-1}},$
@@
$\mathcal{Q}_{EMRI}$ $=0.07m_{\bullet}^{3/2}\left(\frac{125}{m_{\bullet}}\right)^{m_{\bullet}/150}\approx 2.7\left(\frac{m_{\bullet}}{10}\right)^{3/2}.$ (25)
Thus, we get that for $10{M_{\odot}}$ BHs the EMRI rate is $\approx 10^{-8}\ {\rm yr^{-1}}$ and they are enhanced by a factor of $\mathcal{Q}_{EMRI}(10)\approx 2.7$. The scaling $\mathcal{Q}_{EMRI}\propto m_{\bullet}^{3/2}$ stems from two contributions: a factor of $m_{\bullet}^{1/2}$ due to the segregation beyond the sphere of influence (similar to the case of plunges), and a factor of $m_{\bullet}$ from the dependence of the critical radius on the BH mass (Eq. 23).
Furthermore, using Eqs. (21) and (24), we determine the EMRI-to-plunge ratio, a key parameter for predicting the number of sources detectable by LISA (see Babak et al., 2017; Rom et al., 2024)
@@
For $10{M_{\odot}}$ BHs, Eq. (26) predicts an EMRI-to-plunge ratio of $\approx 0.1$. This estimate is consistent with the fiducial value considered by Babak et al. (2017) and is about $1.5$ times lower than the result of Rom et al. (2024), which was based on a two-mass model of the nuclear stellar cluster. The lower ratio arises because, in this work, the $10{M_{\odot}}$ BHs are scattered by heavier $30{M_{\odot}}$ BHs.
@@ 3.4 Stellar EMRI rate
Finally, we consider the case where a main sequence star slowly descends toward the SMBH due to GW emission, i.e., an EMRI of a main sequence star rather than of a BH. In this case, the critical semimajor axis is found by equating the minimal periapsis to the tidal radius (Linial & Sari, 2023), in contrast to $4R_{s}$ as in the BH case. This yields
@@
$\approx 6\times 10^{-4}m^{-0.13}\left(\frac{M_{BH}}{M_{\rm MW}}\right)^{4/3}.$
The resulting main sequence stellar EMRI (sEMRI) rate is
@@
$\approx 2.5\times 10^{-8}m^{-\gamma-0.16}\left(\frac{M_{BH}}{M_{\rm MW}}\right)^{17/12}{\rm yr^{-1}}$
@@
where we assume that stars of mass $m$ are concentrated around $\widetilde{r}_{d}(m)$, as given in Eq. (12). The calculation is further simplified by assuming that for $r\lesssim\widetilde{r}_{d}$ the stellar number density scales as $r^{-3/2}$, neglecting the mass-dependent correction to the power of the density profile (Eq. 1).
The weak $m^{-0.16}$ scaling of the sEMRI rate arises from the critical radius $R_{\star,c}$ (Eq. 27). This reflects the tidal radii variation across different stellar masses, and, hence, implicitly depends on the stellar mass-radius relation. The additional mass-dependent term in the intermediate mass range, $1\lesssim m\lesssim 3$, arises from the segregated location of these stars.
@@
From Eqs. (14) and (28), the sEMRI enhancement factor is
@@
Besides being mHz GW sources, relevant for space-based GW observatories, sEMRIs have recently gained attention as a potential origin of the observed QPEs (Zhao et al., 2021; Linial & Sari, 2023; Lu & Quataert, 2023; Linial & Metzger, 2023; Nicholl et al., 2024). Specifically, Linial & Metzger (2023) associated QPEs with the interaction between a stellar EMRI and a TDE accretion disk.
Given this interpretation, the QPE rate can be estimated by combining the TDE rate (Eq. 21), the sEMRI rate (Eq. 28), and the characteristic EMRI lifetime, $\tau_{GW}\sim m^{-1}{\rm Myr}$ (Linial & Metzger, 2023; Kaur et al., 2024). This yields a characteristic rate $\sim 10^{-6}\rm yr^{-1}$ for one solar-mass stars, roughly compatible with the observational inferred rate (Arcodia et al., 2024b).
@@
We note that beyond the dynamical processes considered in this work, the effects of stellar binaries and collisions should be incorporated for a more detailed analysis of the sEMRIs formation mechanism. For example, binary tidal break-up may be a more efficient formation channel (Linial & Sari, 2023; Lu & Quataert, 2023), while collisions are expected to reduce the abundance of stars on tightly bound orbits, suppressing the sEMRI rate (Sari & Fragione, 2019; Rose et al., 2023; Balberg & Yassur, 2023; Balberg, 2023).
@@ 4 Discussion & Summary
Figure 2: Distributions for stars (yellow lines) and BHs (dashed black lines). Top panel: The number, $N(m)=m{\rm d}N/{\rm d}m$, of stars and BHs within the sphere of influence. The stars follow the stellar PMF (Eq. 4). We show our assumed BH mass distribution (thin gray dashed line, Eq. 6), and its modification due to segregation from beyond the sphere of influence (thick black dashed line, Eq. 10). Second panel: The rates of TDEs, plunges, EMRIs, and sEMRIs. The dashed vertical gray line separates masses that are primarily scattered by stars ($m\lesssim 3$) from those mostly scattered by $30{M_{\odot}}$ BHs. Bottom panel: The enhancement factor $\mathcal{Q}$ for the different transients. Mass segregation within the sphere of influence generally leads to over-representation of transients involving massive stars, resulting in $\mathcal{Q}\approx 9$ for TDEs with $m\gtrsim 3$ (Eq. 18). sEMRIs are suppressed compared to TDEs, but also show over-representation of massive stars. However, for very massive stars ($m\gtrsim 3$), $\mathcal{Q}$ weakly decreases (Eq. 29), reflecting the variation of the tidal radius with stellar mass. For TDEs, $\mathcal{Q}$ sharply increases around $m\approx 3$, as indicated by the gray arrow, since these stars become mostly scattered by BHs. For BHs, segregation beyond the sphere of influence takes place as well, leading to an enhancement factor that increases with mass. For EMRIs, $\mathcal{Q}$ rises more steeply due to the additional dependence of the EMRI critical radius on the BH mass (Eq. 23).
We study the distributions of stars and BHs in nuclear stellar clusters surrounding SMBHs. We highlight the impact of mass segregation on the rates of TDEs, plunges, and EMRIs.
@@
Furthermore, we assume that $30{M_{\odot}}$ BHs are sufficiently abundant to dominate over all other stellar-mass BHs. This is inspired by the emerging BH mass spectrum, inferred from the observations of the LVK collaboration (Abbott et al., 2023). Given this assumption, the specific details of the BH distribution – taken here as a power-law toy model with tails extending to both low and high mass ends – affect the quantitative rate estimates but do not alter the qualitative trends discussed in this work.
We present a simple derivation of the segregated steady-state stellar distribution, in agreement with a previous analytic calculation of Linial & Sari (2022). Additionally, we schematically outline the BH distribution and its impact on the stellar distribution, consistently accounting for the mutual interactions between BHs and stars.
Accounting for the stellar-mass BHs, we show that their number within the sphere of influence increases by a factor of $\approx 3$ relative to the expected number based on the stellar IMF, as their relaxation time is shorter than the age of the galaxy, even outside the sphere of influence, allowing them to sink inwards from greater distances.
We further estimate the EMRI-to-plunge ratio and show that scattering by more massive BHs reduces the EMRI rate. This is because the scattering by heavier BHs shifts the EMRIs critical radius toward smaller distances, where the BH population is less abundant. Nonetheless, the segregated BH distribution leads to an over-representation of EMRIs and plunges involving massive stellar-mass BHs, with enhancement factor scaling as $\mathcal{Q}_{EMRI}\propto m_{\bullet}^{3/2}$ and $\mathcal{Q}_{Plunge}\propto m_{\bullet}^{1/2}$, respectively. These results are of particular interest, as they influence the expected number of EMRIs to be measured by LISA (Babak et al., 2017; Amaro-Seoane et al., 2023; Rom et al., 2024).
@@
Finally, we estimate the sEMRI rate, showing that it favors low-mass stars ($m\lesssim 1$). We find that the sEMRI rate is reduced compared to its expected rate from single-mass models (e.g., Linial & Sari, 2023), yet remains marginally consistent with the observationally inferred QPE rate (Arcodia et al., 2024b). However, a more detailed analysis, including the effects of stellar binaries and collisions (Rose et al., 2023; Balberg & Yassur, 2023; Balberg, 2023), is necessary for accurately determining the sEMRI rate.

### v2

@@ 3.1 TDE rate
The majority of disrupted stars with masses in the range $1\lesssim m\lesssim 3$ originate from the vicinity of $r_{d}(m)$, as defined by Equation (9), where they dominate the scatterings. In this region, the relaxation timescale (Equation 8) is
@@
At smaller distances, both their number and the two-body scattering flux decrease as a power law, and at larger distances their number is exponentially suppressed (Linial & Sari, 2022).
Therefore, using Equations (4), (13), and (15), the mass-dependent tidal disruption rate is given by
@@
[…]
Consequentially, the enhancement factor (defined in Equation 14) is
@@
Given that $\gamma\approx 5$ for this stellar-mass range (Equation 5), the TDE rate scales roughly as ${\rm d}\Gamma/{\rm d}m\propto m^{-4}$ and the enhancement factor is $\mathcal{Q}_{TDE}\simeq m$. Notably, the TDE rate, as well as the rates of EMRIs and plunges (calculated below), scales with the SMBH mass as $\Gamma\propto M_{BH}^{-1/4}$ (in agreement with previous results, e.g., Hopman & Alexander, 2005; Kochanek, 2016; Yao et al., 2023).
As discussed in Section 2.4, the presence of a BH population modifies the cusp within $R_{\bullet}$. The transition of massive stars ($m\gtrsim\widetilde{m}$) from being primarily self-scattered to being scattered by BHs results in a sharp increase in the enhancement factor [^5] This increase originates from the shorter relaxation timescale when scattered by the more massive BHs compared to the stars, $\tau_{\rm rlx}(\widetilde{m})/\tau_{\rm rlx}(\widetilde{m}_{\bullet})\approx 3$, as evident from Equations (15) and (20). around $R_{\bullet}$, $\delta\mathcal{Q}\approx\left(\widetilde{m}_{\bullet}/\widetilde{m}\right)^{1/2}\approx 3$. Therefore
[…]
@@
The enhancement factor of stellar disruptions is illustrated in Figure (2) and can be summarized as follows:
@@ 3.2 Plunge rate
The stellar-mass BHs are predominantly scattered by the $\widetilde{m}_{\bullet}\approx 30$ BHs near $R_{\bullet}$. Therefore, the relevant scattering timescale (Equation 8) for the plunge rate is given by
@@
[…]
@@
This scaling reflects the flattening of the BH mass spectrum caused by the segregation into the sphere of influence (see Section 2.4). It applies to both lighter and heavier BHs [^6] For massive stellar-mass BHs, with $m_{\bullet}\gtrsim 125$, The critical radius $R_{c}$ (Equation 23), which differentiate between plunges and EMRIs, exceeds $R_{\bullet}$ (Equation 11), the radius around which the BHs concentrate. Therefore, such BHs are more likely to form EMRIs than undergo a plunge., since they are mostly scattered by BHs with mass $\widetilde{m}_{\bullet}$ around $R_{\bullet}$, leading to $\left.{\rm d}\Gamma/{\rm d}m\right|_{Plunge}\propto m_{\bullet}^{-\widetilde{\gamma}_{\bullet}}$.
@@ 3.3 EMRI rate
[…]
@@
[…]
$\quad\times\left(\frac{M_{BH}}{M_{\rm MW}}\right)^{-1/4}{\rm yr^{-1}},$
@@
[…]
Thus, we get that for $10{M_{\odot}}$ BHs the EMRI rate is $\approx 10^{-7}\ {\rm yr^{-1}}$ and they are enhanced by a factor of $\mathcal{Q}_{EMRI}(10)\approx 2.7$. The scaling $\mathcal{Q}_{EMRI}\propto m_{\bullet}^{3/2}$ stems from two contributions: a factor of $m_{\bullet}^{1/2}$ due to the segregation beyond the sphere of influence (similar to the case of plunges), and a factor of $m_{\bullet}$ from the dependence of the critical radius on the BH mass (Equation 23).
Furthermore, using Equations (21) and (24), we determine the EMRI-to-plunge ratio, a key parameter for predicting the number of sources detectable by LISA (see Babak et al., 2017; Rom et al., 2024)
@@
For $10{M_{\odot}}$ BHs, Equation (26) predicts an EMRI-to-plunge ratio of $\approx 0.1$. This estimate is consistent with the fiducial value considered by Babak et al. (2017) and is about $1.5$ times lower than the result of Rom et al. (2024), which was based on a two-mass model of the nuclear stellar cluster. The lower ratio arises because, in this work, the $10{M_{\odot}}$ BHs are scattered by heavier $30{M_{\odot}}$ BHs.
@@ 3.4 Stellar EMRI rate
Finally, we consider the case where a main-sequence star slowly descends toward the SMBH due to GW emission, i.e., an EMRI of a main-sequence star rather than of a BH. In this case, the critical semimajor axis is found by equating the minimal periapsis to the tidal radius (Linial & Sari, 2023), in contrast to $4R_{s}$ as in the BH case. This yields
@@
[…]
The resulting main-sequence stellar EMRI (sEMRI) rate is
@@
[…]
where we assume that stars of mass $m$ are concentrated around $\widetilde{r}_{d}(m)$, as given in Equation (12). The calculation is further simplified by assuming that for $r\lesssim\widetilde{r}_{d}$ the stellar number density scales as $r^{-3/2}$, neglecting the mass-dependent correction to the power of the density profile (Equation 3).
The weak $m^{-0.16}$ scaling of the sEMRI rate arises from the critical radius $R_{\star,c}$ (Equation 27). This reflects the tidal radii variation across different stellar masses, and, hence, implicitly depends on the stellar-mass-radius relation. The additional mass-dependent term in the intermediate mass range, $1\lesssim m\lesssim 3$, arises from the segregated location of these stars.
@@
From Equations (14) and (28), the sEMRI enhancement factor is
@@
Besides being mHz GW sources, relevant for space-based GW observatories, sEMRIs have recently gained attention as a potential origin of the observed QPEs (Zhao et al., 2021; Linial & Sari, 2023; Lu & Quataert, 2023; Linial & Metzger, 2023; Nicholl et al., 2024). Specifically, Linial & Metzger (2023) associated QPEs with the interaction between a sEMRI and a TDE accretion disk.
Given this interpretation, the QPE rate can be estimated by combining the TDE rate (Equation 21), the sEMRI rate (Equation 28), and the characteristic EMRI lifetime, $\tau_{GW}\sim m^{-1}{\rm Myr}$ (Linial & Metzger, 2023; Kaur et al., 2024). This yields a characteristic rate $\sim 10^{-6}\rm yr^{-1}$ for one solar-mass stars, roughly compatible with the observational inferred rate (Arcodia et al., 2024b).
@@
We note that beyond the dynamical processes considered in this work, the effects of stellar binaries and collisions should be incorporated for a full analysis of the sEMRI formation mechanism. For example, binary tidal break-up may be a more efficient formation channel (Linial & Sari, 2023; Lu & Quataert, 2023), while collisions are expected to reduce the abundance of stars on tightly bound orbits, suppressing the sEMRI rate (Sari & Fragione, 2019; Rose et al., 2023; Balberg & Yassur, 2023; Balberg, 2024).
@@ 4 Discussion & Summary
Figure 2: Distributions for stars (yellow lines) and BHs (dashed black lines). Top panel: The number, $N(m)=m{\rm d}N/{\rm d}m$, of stars and BHs within the sphere of influence. The stars follow the stellar PMF (Equation 4). We show our assumed BH mass distribution (thin gray dashed line, Equation 6), and its modification due to segregation from beyond the sphere of influence (thick black dashed line, Equation 10). Second panel: The rates of TDEs, plunges, EMRIs, and stellar EMRIs (sEMRIs). The dashed vertical gray line separates masses that are primarily scattered by stars ($m\lesssim 3$) from those mostly scattered by $30{M_{\odot}}$ BHs. Bottom panel: The enhancement factor $\mathcal{Q}$ for the different transients. Mass segregation within the sphere of influence generally leads to over-representation of transients involving massive stars, resulting in $\mathcal{Q}\approx 9$ for TDEs with $m\gtrsim 3$ (Equation 18). The sEMRIs are suppressed compared to TDEs, but also show over-representation of massive stars. However, for very massive stars ($m\gtrsim 3$), $\mathcal{Q}$ weakly decreases (Equation 29), reflecting the variation of the tidal radius with stellar mass. For TDEs, $\mathcal{Q}$ sharply increases around $m\approx 3$, as indicated by the gray arrow, since these stars become mostly scattered by BHs. For BHs, segregation beyond the sphere of influence takes place as well, leading to an enhancement factor that increases with mass. For EMRIs, $\mathcal{Q}$ rises more steeply due to the additional dependence of the EMRI critical radius on the BH mass (Equation 23).
We study the distributions of stars and BHs in nuclear stellar clusters surrounding SMBHs. We analytically study the impact of two-body scattering and the resulting mass segregation, and self-consistently derive the steady-state distributions for a population of stars and stellar-mass BHs. Based on these distributions, we estimate the formation rate of various transients and show that mass segregation generally enhances the fraction of transients involving massive stars and stellar-mass BHs relative to their overall abundances.
@@
Additionally, we assume that $30{M_{\odot}}$ BHs are sufficiently abundant to dominate over all other stellar-mass BHs. This is inspired by the emerging BH mass spectrum, inferred from the observations of the LVK collaboration (Abbott et al., 2023). Given this assumption, the specific details of the BH distribution - taken here as a power-law toy model with tails extending to both low and high mass ends - slightly affect the quantitative rate estimates but do not alter the qualitative trends discussed in this work. Furthermore, in the case that the $10{M_{\odot}}$ BHs are the dominant scatterers, with the overall BH number fraction held fixed, the scattering timescale and the characteristic distances for EMRI and sEMRI formation would change only by factors of order-unity. Notably, while the formation rates depend on the shape of the BH mass function, as they scale linearly with the number of BHs at each mass, this dependence is factored out in the enhancement factor $Q$. The enhancement trends, pointed out in our analysis, reflect dynamical effects and are largely insensitive to the specific form of the mass function.
We present a simple derivation of the segregated steady-state distributions for both the stars and the stellar-mass BHs, accounting for their mutual interactions. These distributions are obtained using our generalized formulation of the constant energy flux condition, which enables the derivation of several new results.
We show that the number of BHs within the sphere of influence increases by a factor of $\approx 3$ relative to the expectations based on the stellar IMF, since their relaxation time is shorter than the age of the galaxy, even beyond the sphere of influence, enabling them to sink inward from larger radii. Furthermore, we find that BHs dominate the scattering within $R_{\bullet}\sim 0.1R_{h}$. Therefore, in this inner region, they form a BW cusp. This prediction is consistent with the findings of several numerical studies (e.g., Freitag et al., 2006; Panamarev et al., 2019; Broggi et al., 2022; Balberg, 2024).
Taking into account the formation of transients, we estimate the EMRI-to-plunge ratio and show that scattering by more massive BHs reduces the EMRI rate. This is because the scattering by heavier BHs shifts the EMRIs critical radius toward smaller distances, where the BH population is less abundant. Nonetheless, the segregated BH distribution leads to an over-representation of EMRIs and plunges involving massive stellar-mass BHs, with enhancement factor scaling as $\mathcal{Q}_{EMRI}\propto m_{\bullet}^{3/2}$ and $\mathcal{Q}_{Plunge}\propto m_{\bullet}^{1/2}$, respectively. These results are of particular interest because they influence the expected number of EMRIs to be measured by LISA (Babak et al., 2017; Amaro-Seoane et al., 2023; Rom et al., 2024).
@@
Finally, we estimate the sEMRI rate, showing that it favors low-mass stars ($m\lesssim 1$). We find that the sEMRI rate is reduced compared to its expected rate from single-mass models (e.g., Linial & Sari, 2023), yet remains marginally consistent with the observationally inferred QPE rate (Arcodia et al., 2024b). However, a more detailed analysis, including the effects of stellar binaries and collisions (Rose et al., 2023; Balberg & Yassur, 2023; Balberg, 2024), is necessary for accurately determining the sEMRI rate.